A very long string (linear density 0.7 kg/m ) is stretched with a tension of 85 N . One end of the string oscillates up and down with an amplitude of 7 cm and a period of 0.35 s . What is the wavelength of the waves created in the string?
Answer #1

Solution) Linear density , U = 0.7 kg/m

Tension , T = 85 N

Amplitude , A = 7 cm = 0.07 m

Period , t = 0.35 s

Wavelength, Lambda = ?

V = (T/U)^(1/2)

V = (85/0.7)^(1/2)

V = 11.02 m/s

V = f(Lambda)

Lambda = (V/f)

f = 1/t

f = 1/(0.35)

f = 2.86 Hz

Lambda = (11.02/2.86)

Lambda = 3.85 m

Wavelength = 3.85 m